This homework will be due on Wednesday Feb 19th in discussion. Question 1. A cell of some bacteria divides into two cells eveCan you answer questions 2-6 please? Thanks
Answer #1

Solution Orthogonal trajectories of the family of curve 92=6cm Given that y ²6cn - 1) Differentiate equ (1) wisit n we get 29Solution Given that y-69+99=0 Given equation can be written an. (D2-6D + 9 )y=0 Auxiliary equation is m26m +9 = 0 {m=Die masolution 5 Given that 94(1) +9(n)-6y=0 This can be written as (D4 + D²-6)y=0 Auxilary equation is given by mitm?_6=0 Here m-